Jse appropriate algebra and Theorem 7.2.1 to 45 52 +9 Talk to a Tutor Read it Need Help? Submit Answer
Answer #1


6 S +9 s +9 (5 +7 (45-61 Given that - 13+9) 45-6 4s Consider s +9 4s - 4s 52 +9 52 +9 4s - -6 42- s? +9 45-6 1-2 = 42- 2 }= b